OpenStudy (arianawo2):

A MEDAL!! There are 30 homes in Neighborhood A. Each year, the number of homes increases by 20%. Just down the road, Neighborhood B has 45 homes. Each year, 3 new homes are built in Neighborhood B. Part A: Write functions to represent the number of homes in Neighborhood A and Neighborhood B throughout the years. Part B: How many homes does Neighborhood A have after 5 years? How many does Neighborhood B have after the same number of years? Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically

OpenStudy (arianawo2):

@soprano.h.d0816 here it is

OpenStudy (soprano.h.d0816):

Well. I am confusing myself on this one. Lol

OpenStudy (arianawo2):

I know right, but I just need help with part B and C because I have this Part A: Neighborhood A f(x)=30∗0.2x Neighborhood B (x)=45+3x

OpenStudy (soprano.h.d0816):

For part B, you'd plug in 5 for x

OpenStudy (arianawo2):

Yah but it would be like this? f(x)=30(1.2)^5 30(2.4) 72 f(x)=45+36^5=288

OpenStudy (soprano.h.d0816):

Where did you get 1.2 from?

OpenStudy (arianawo2):

So, is not 1.2? is 0.2? because is increasing I think

OpenStudy (soprano.h.d0816):

I'm so confused now. I'm sorry

OpenStudy (arianawo2):

Don't worry, thanks for trying tho, have a great day! :)

OpenStudy (soprano.h.d0816):

You too hun(:

OpenStudy (arianawo2):

@KyleWells17 this is my question

OpenStudy (anonymous):

Oh damn.

OpenStudy (arianawo2):

@Badsheep

OpenStudy (arianawo2):

@AloneS

OpenStudy (arianawo2):

I already tried, no one has the answer :(